# The anonymization problem in social networks

### v1

@@ Abstract
In this paper we introduce a general version of the anonymization problem in social networks, in which the goal is to maximize the number of anonymous nodes by altering a given graph. We define three variants of this optimization problem, being full, partial and budgeted anonymization. In each, the objective is to maximize the number of $k$-anonymous nodes, i.e., nodes for which there are at least $k-1$ equivalent nodes, according to a particular anonymity measure of structural node equivalence. We propose six new heuristic algorithms for solving the anonymization problem which we implement into the reusable ANO-NET computational framework. As a baseline, we use an edge sampling method introduced in previous work. Experiments on both graph models and 17 real-world network datasets result in three empirical findings. First, we demonstrate that edge deletion is the most effective graph alteration operation. Second, we compare four commonly used anonymity measures from the literature and highlight how the choice of anonymity measure has a tremendous effect on both the achieved anonymity as well as the difficulty of solving the anonymization problem. Third, we find that the proposed algorithms that preferentially delete edges with a larger effect on nodes at a structurally unique position consistently outperform heuristics solely based on network structure. With similar runtimes, our algorithms retain on average 17 times more edges, ensuring higher data utility after full anonymization. In the budgeted variant, they achieve 4.4 times more anonymous nodes than the baseline. This work lays important foundations for future development of algorithms for anonymizing social networks.
@@ 1 Introduction
The field of social network analysis, also referred to as network science, studies interactions between people for various applications and downstream tasks, including modeling disease spread [1], maximizing influence spread [2], spotting anomalies [3] and finding communities [4]. To conduct such research, it is desirable to have large real-world networks about people that are as realistic as possible. However, various studies found that in these networks people can still be identified by using structural information, even after removing unique identifiers, i.e., pseudonymization [5, 6, 7]. As a result, sharing or publishing such networks can lead to a breach in privacy of the people represented.
@@
To solve this problem, sometimes also referred to as identity anonymization [8], various works introduced methods to share or publish networks in a privacy aware manner, including clustering [9, 10, 11, 12], differential privacy [13, 14, 15, 16, 17, 18, 19]nd $k$-anonymity [20, 21, 6, 22]. Each of these approaches anonymizes the data in a different way. For clustering, an intermediate version of the network is shared, where nodes are clustered together into supernodes and superedges in a way that anonymity is assured. While this representation can preserve global properties, local structure can not be preserved. In turn, in differential privacy, users can run a query to obtain modestly perturbed information about the network. Even though several downstream network analysis tasks can be addressed using differential privacy [23, 24], there are a number of practical settings in which a social network dataset must first be published or shared with a third party before it can be analyzed. This can be because two parties work together, e.g., where one collects and the other analyzes the data. Therefore, in this work we focus on $k$-anonymity methods, which aim to create an altered anonymized version of the original full network dataset.
@@
In $k$-anonymity approaches, the anonymity of a network as a whole is determined by splitting the nodes into equivalence classes based on a chosen notion for node equivalence, which we refer to as the anonymity measure. The choice of measure essentially determines the attacker scenario against which one protects [25]. Examples are the degree based attack [20, 26, 27, 28, 29, 30, 8] or the neighborhood attack, accounting for the exact structure of a node’s neighborhood [6, 31, 32, 33, 34, 7]. A node is $k$-anonymous if it is structurally equivalent to $k-1$ other nodes, according to the used measure for anonymity. For example, if a node is $k$-anonymous with respect to the degree measure, there are $k-1$ nodes with the same degree.
Previous work focuses on the situation where $k=2$, allowing the anonymity of the network to be measured as uniqueness: the fraction of nodes at a structurally unique network position [6, 25]. Given a network and an anonymity measure, the aim of network anonymization is then to alter the network in order to make more nodes anonymous, i.e., fewer nodes unique.
In network anonymization there is a trade-off between anonymity and utility. Utility can be defined in many ways and is dependent on the applications the anonymized network should be used for. Since in general altering fewer edges leads to higher utility in the form of preservation of structural network features and centrality [35], in this paper we focus on altering as few edges as possible.
@@
The time complexity of anonymizing a network with as few alterations as possible depends on the chosen anonymity measure. For example, [8] showed that anonymizing for the degree measure can be done in quadratic time in terms of the number of nodes. However,  [32], which uses as measure the exact structure of the 1-neighborhood of a considered node, showed that anonymization is NP-complete. Furthermore, the anonymization problem is not monotonic nor submodular, elegant properties that enabled the development of efficient methods for problems such as influence maximization [36]. As a result, existing works on anonymization of networks center around (meta)heuristic algorithms. A number of these algorithms are designed for a specific measure [20, 32], while those that are measure-agnostic typically perform alterations randomly [6, 37]. In addition, some works used genetic algorithms [34, 26]. The algorithms vary in the perturbation operations used, and are sometimes tailored to a specific type of network [38].
@@
In this paper, we introduce a general version of the anonymization problem in the form of three meaningful variants, being 1) full, 2) partial and 3) budgeted anonymization. Together with this paper, we release a computational framework for anonymizing networks called ANO-NET that allows researchers to implement their anonymization algorithms and apply these to the different variants of the anonymization problem. The framework is accompanied by a representative real-world benchmark dataset. Additionally, we introduce six new heuristic measure-agnostic anonymization algorithms. We focus on edge deletion as our empirical results show this is the most effective operation and because this was found to be less damaging to data utility than edge addition [35]. We anonymize against four different anonymity measures to gain insight into the effect of the chosen measure on the anonymization process. Our results show the effectiveness of our heuristic based algorithms, which target specific edges based on network structure or uniqueness of the nodes.
As existing algorithms use different sets of operations or are not measure-agnostic, we use edge sampling (es[6] as our baseline. On average, our best heuristic algorithm makes 4.4 times more nodes anonymous for the budgeted variant, and in full anonymization 17 times more edges are preserved than using the es baseline.
@@
  • We propose six new measure-agnostic heuristic based algorithms for anonymization.
@@
It is worth noting that while introduced for social networks, privacy can also be important for other types of networks, such as software design networks [38] or networks of companies [39]. Hence, in our set of real-world networks we include different categories of network datasets besides social networks.
The remainder of this paper is structured as follows. First, Section 2 summarizes related work on the problem. Section 3 covers preliminaries about networks and anonymization. In Section 4, we introduce the anonymization problem and its variants. Next, in Section 5, we outline our framework for the anonymization process and present the proposed heuristic anonymization algorithms. We discuss experiments and results in Section 6. Lastly, we give a summary and directions for future work in Section 7.
@@ 2 Related work
Usually, works introduced in the literature aim to anonymize all nodes in the network, the first variant of the problem. They often focus on a specific anonymity measure, and based on this create an anonymization algorithm. The time complexity of the anonymization problem then also depends on the chosen measure. Below, we summarize for the different measures the complexity of anonymizing the network using as few alterations as possible, as well as the anonymization algorithms introduced in these works.
@@
Degree. The most simple anonymity measure is the degree of the node [20, 26, 27, 28, 29, 30, 8]. As a result, exact anonymization algorithms have been introduced which require the smallest number of alterations and run in quadratic time in terms of the number of nodes [20, 30, 8]. With the exception of [26], which introduces a genetic algorithm for anonymization, most of these works introduce an anonymization algorithm specific to the degree measure, such as the UMGA [30] or KDA algorithm [8].
@@
Neighborhood. More complex are the measures accounting for the exact structure of the ego network, i.e., 1-neighborhood [6, 34, 32], or as a parameterized version, the $d$-neighborhood structure [40]. When using this measure, the anonymization problem has been shown to be NP-complete [32]. As a result, no exact algorithms have been introduced to solve this problem. Works have introduced greedy algorithms aiming to change neighborhoods of nodes into that of similar nodes [32], a genetic algorithm [34] and an algorithm based on edge sampling [6].
@@
Graph invariants. The work of [25] included two new measures based on graph invariants, being the number of nodes and edges in the $d$-neighborhood of a considered node and its degree distribution. Complexity or anonymization approaches were not discussed.
@@
VRQ. The notion of Vertex Refinement Queries (VRQ) was introduced in the work of [37, 21, 41], accounting for the degree of nodes that are at most at distance $d$ of the considered node. To anonymize, algorithms using random rewiring [37] and alterations based on finding similar neighborhoods [41] were introduced. The work in [42] proves that anonymization using this maesure is NP-complete for the subset variant which aims to anonymize a subset of nodes.
@@
In this paper we add to existing literature by introducing and investigating three meaningful variants of the anonymization problem. Additionally, we research the effect of different anonymity measures on the anonymization process and introduce, propose and compare seven measure-agnostic anonymization algorithms.
@@ 3 Preliminaries
In this section we introduce concepts and terminology used throughout this paper. This includes general concepts for networks and anonymity, and the used anonymity measures.
3.1 Networks and anonymization
We define a network as an undirected graph $G=(V,E)$ containing a set of nodes $V$ and set of edges $\{v,w\}\in E$ between nodes $v,w\in V$. The degree of a node equals the number of connections it has: $degree(v)=|\{w:\{v,w\}\in E\}|$. The distribution of node degrees in a real-world network typically resembles a powerlaw distribution with many low degree nodes, and few high degree nodes [43].
Nodes in networks often tend to cluster. This tendency is captured for a node $v\in V$ by the clustering coefficient which equals the number of triangles it is part of, divided by the maximum number of triangles it could have: $\frac{1}{2}degree(v)(degree(v)-1)$. The clustering coefficient of a network equals the average clustering coefficient of the nodes in the network. The tendency of nodes to connect to nodes with (dis)similar degrees is measured by assortativity. A positive value implies the nodes connect to similar nodes and a negative value implies the node tends to connect to dissimilar nodes. The distance between two nodes, $dist(v,w)$, equals the number of edges that needs to be traversed to move from one node to the other. It follows that $dist(v,v)=0$ and if there is no path between the nodes we define $dist(v,w)=\infty$.
We define the $d$-neighborhood of a node $N_{d}(v)=(V_{N_{d}(v)},E_{N_{d}(v)})$ as the subgraph consisting of the set of nodes that are at most at distance $d$ of the considered node $v$, and the set of all edges between these nodes. Two $d$-neighborhoods are structurally indistinguishable if they are isomorphic, i.e, the node sets are connected by a bijection that preserves all connections. We can determine if two $d$-neighborhoods are isomorphic by comparing their canonical labeling [44]. This is a label assigned by a function $\mathcal{C}()$ such that two graphs have the same label value only if they are isomorphic.
@@ 3.2 Anonymity in networks
For a graph and an anonymity measure $M$, the fact that two nodes $v,w\in V$ are equivalent is denoted $v\cong_{M}w$, if $M(v)=M(w)$. Two such nodes are said to be in the same equivalence class, denoted $eq_{M}(v)$ for a node $v$. A node is $k$-anonymous if $|eq_{M}(v)|=k$ and $P_{M}$ denotes the partitioning of nodes into equivalence classes using measure $M$. When choosing $k=2$, the anonymity of a graph can be summarized using the metric of uniqueness:
$uniqueness(G)=\frac{|\{v\in V:|eq_{M}(v)|=1\}|}{|V|}$
@@ 3.2.1 Alteration operations
When anonymizing a graph, the aim is to make as many nodes as possible $k$-anonymous given a chosen measure for anonymity. While in the majority of this paper we focus on edge deletion, we experiment with the following edge alteration operations:
  • Deletion: $E\leftarrow E\setminus\{\{v,w\}\},s.t.\{v,w\}\in E$.
  • Addition: $E\leftarrow E\cup\{\{v,w\}\},s.t.\{v,w\}\notin E$.
  • Rewiring: $E\leftarrow E\setminus\{\{v,w\},\{v^{\prime},w^{\prime}\}\}\cup\{\{v,w^{\prime}\},\{v^{\prime},w\}\}$
@@
It should be noted that in many downstream tasks where the objective is to classify a particular aspect of a node, or to cluster nodes, operations that affect the nodes are often not desired. Therefore, we do not consider node deletion or addition.
Table 1: Anonymity measures and their abbreviation (left two columns), definition given a graph $G=(V,E)$, considered node $v$ and distance $d$ (third column), reach of the measure, and node set affected when altering edge $\{v,w\}$ (last two columns).
Measure Abbreviation $M(v,d)$ Reach $A_{M}(\{v,w\})$

### v2

@@ Abstract
In this paper we introduce a general version of the anonymization problem in social networks, in which the goal is to maximize the number of anonymous nodes by altering a given graph. We define three variants of this optimization problem being full, partial and budgeted anonymization. In each, the objective is to maximize the number of $k$-anonymous nodes, i.e., nodes for which there are at least $k-1$ equivalent nodes, according to a particular anonymity measure of structural node equivalence. We propose four new heuristic algorithms for solving the anonymization problem which we implement into a reusable computational framework. As a baseline, we use an edge sampling method introduced in previous work. Experiments on both graph models and 19 real-world network datasets result in three empirical findings. First, we demonstrate that edge deletion is the most effective graph alteration operation. Second, we compare four commonly used anonymity measures from the literature and highlight how the choice of anonymity measure has a tremendous effect on both the initial anonymity as well as the difficulty of solving the anonymization problem. Third, we find that the proposed algorithm that preferentially deletes edges with a larger effect on nodes at a structurally unique position consistently outperforms heuristics solely based on network structure. Our best performing algorithm retains on average 14 times more edges in full anonymization, and overall ensures a better trade-off between anonymity and data utility. In the budgeted variant, it achieves 4.8 times more anonymous nodes than the baseline. This work lays foundations for future development of algorithms for anonymizing social networks.
keywords:
Complex networks, $k$-Anonymity, Privacy, Algorithms
@@ 1 Introduction
The field of social network analysis, also known as network science, studies interactions between people for various applications and downstream tasks, including modeling disease spread [1], maximizing influence spread [2], spotting anomalies [3] and finding communities [4, 5]. To conduct such research, it is desirable to have large real-world networks of people that are as realistic as possible. However, various studies found that people can still be identified in these networks using structural information, even after removing unique identifiers, i.e., pseudonymization [6, 7, 8]. As a result, sharing or publishing such networks can lead to a breach in the privacy of the people represented.
@@
To solve this problem, sometimes also referred to as identity anonymization [9], various works introduced methods to share or publish networks in a privacy aware manner including clustering [10, 11, 12, 13], differential privacy [14, 15, 16, 17, 18, 19, 20] and $k$-anonymity [21, 22, 7, 23]. Each of these approaches anonymizes the data in a different way. For clustering, an intermediate version of the network is shared, where nodes are clustered together into supernodes and superedges in a way that anonymity is ensured. While this representation can preserve global network properties, the local structure can not be preserved. In turn, in differential privacy, users can run a query to obtain perturbed information about the network. After each query, progressively more noise should be added to ensure privacy. Even though several downstream network analysis tasks can be addressed using differential privacy [24, 25], there are a number of practical real-world settings in which a social network dataset must first be published or shared with a third party before it can be analyzed. This can be because two parties collaborate, for example, one collects and the other analyzes the data. Therefore, in this work, we focus on the $k$-anonymity methods, which aim to create an altered anonymized version of the original full network dataset.
@@
In $k$-anonymity approaches, the anonymity of a network as a whole is determined by partitioning the nodes into equivalence classes based on a chosen notion for node equivalence, which we refer to as the anonymity measure. A node is $k$-anonymous if it is structurally equivalent to $k-1$ other nodes, according to the measure used for anonymity. For example, if a node is $k$-anonymous with respect to the degree measure, there are $k-1$ nodes with the same degree. The choice of measure essentially determines the attacker scenario against which one protects [26]. Examples are the degree based attack [21, 27, 28, 29, 30, 31, 9] or the neighborhood attack, accounting for the exact structure of a node’s neighborhood [7, 32, 33, 34, 35, 8]. While in this manuscript we focus on unlabeled and undirected graphs, the anonymity measures can be extended to account for additional attacker knowledge such as node labels [33, 34, 36, 37], edge labels [38] or edge weights [39].
Previous work focuses on the situation where $k=2$, where the anonymity of the network is measured as uniqueness: the fraction of nodes at a structurally unique network position [7, 26]. Given a network and an anonymity measure, the aim of network anonymization is then to decrease the uniqueness to a desired level, by means of network alterations, while ensuring that the resulting network data remains suitable for its intended purposes. Thus ultimately, in network anonymization there is a trade-off between anonymity and data utility. How utility is defined depends on the applications for which the anonymized network should be used. To measure utility, we distinguish between preserving structural network properties, and performance on downstream tasks. This distinction is similar to work in statistical disclosure control [40] that distinguishes between general utility, the distribution of the data, hence describing the data itself, and specific utility, concerning results of analyzing the data. Previous work has found that altering fewer edges leads to higher utility in the form of preservation of structural network features and centrality ranking of nodes [41]. Hence, in this paper we focus on altering an as small number of edges as possible.
@@
The time complexity of anonymizing a network with as few alterations as possible depends on the chosen anonymity measure. For example, [9] showed that anonymization for the degree measure can be done in quadratic time in terms of the number of nodes. However,  [33], which uses as measure the exact structure of the 1-neighborhood of a considered node, showed that anonymization is NP-complete. Furthermore, the objective function in the anonymization problem is neither monotonic nor submodular, elegant properties that enabled the development of efficient methods for common network analysis problems such as influence maximization [42]. As a result, existing work on the anonymization of networks focuses on (meta)heuristic algorithms. Several of these algorithms are designed for a specific measure [21, 33], while those that are measure-agnostic typically perform alterations randomly [7, 43]. In addition, some works used genetic algorithms [35, 27]. The algorithms vary in the alteration operations used, and are sometimes tailored to a specific type of network [44].
@@
In this paper, we introduce a general version of the anonymization problem in the form of three meaningful variants, being 1) full, 2) partial and 3) budgeted anonymization. By means of experiments, we investigate different aspects of the problem. We focus on edge deletion as alteration operation, because our empirical results show that this is the most effective operation when randomly applied to a network. We introduce four new heuristic measure-agnostic anonymization algorithms, which we compare against a baseline edge sampling algorithm [7]. We anonymize against four different anonymity measures to gain insight into the effect of the chosen measure on the anonymization process, from which we choose one measure to focus on. Then, we compare the performance of the proposed algorithms on the three variants of the problem and assess the utility of (partially) anonymized networks.
Our results show the effectiveness of our heuristic based algorithms, which target specific edges based on network structure or uniqueness of the nodes. On average, compared to the baseline, our best heuristic algorithm makes 4.8 times more nodes anonymous for the budgeted variant, and in full anonymization 13.9 times more edges are preserved. Additionally, our best heuristic algorithm preserves utility at least as well as the baseline approach, given the same number of edge deletions, and results in a better trade-off between anonymity and utility. Together with this paper, we release a computational framework for anonymizing networks called ANO-NET [^1] https://github.com/RacheldeJong/ANONET that allows researchers to implement their anonymization algorithms and apply them to the different variants of the anonymization problem. The framework is accompanied by a representative real-world benchmark dataset.
@@
  • We propose four new measure-agnostic heuristic based algorithms for anonymization.
@@
  • We assess the utility of the (partially) anonymized graphs and investigate the anonymity versus utility trade-off achieved by different algorithms.
It is worth noting that while introduced for social networks, privacy can also be important for other types of networks, such as software design networks [44] or networks of companies [45]. Hence, in our set of real-world networks, we include different categories of network datasets besides social networks.
The remainder of this paper is structured as follows. First, Section 2 summarizes related work on the problem. Section 3 covers preliminaries on networks and anonymization. In Section 4, we introduce the anonymization problem and its variants. Next, in Section 5, we outline our framework for the anonymization process and present the proposed heuristic anonymization algorithms. We discuss experiments and results in Section 6. Lastly, we give a summary and directions for future work in Section 7.
@@ 2 Related work
Works on the anonymization problem introduced in the literature often aim to anonymize all nodes in the network, the first variant of the problem. They typically focus on a specific anonymity measure, and based on this create an anonymization algorithm [21, 33]. The chosen measure also affects the complexity of the problem. For example, when the chosen measure is the degree of the nodes and all edge operations are allowed, the problem could be solved in $\mathcal{O}(n^{2})$, where $n$ is the number of nodes in the network [21]. However, for a more complex measure, such as the exact structure surrounding a node, the problem is NP-hard [32]. The search space for possible solutions is tremendously large, consisting of all subsets of existing edges, or even all subsets of possible edges, when allowing for edge addition. Therefore, exact approaches are not suitable and there is a clear need for approximation algorithms to handle larger graph inputs. Below we give an overview of related work and anonymization algorithms introduced categorized by the anonymity measure used in the work.
@@
Degree. The most simple anonymity measure is the degree of the node [21, 27, 28, 29, 30, 31, 9]. As a result, exact anonymization algorithms have been introduced that require a minimum number of alterations [21, 31, 9]. A different approach is introduced in [27], which introduces a genetic algorithm for anonymization. Most existing work introduces an anonymization algorithm specific to the degree measure, such as the UMGA [31] or KDA algorithm [9].
@@
Neighborhood. Anonymity measures that account for more complex information are based on the exact structure of the ego network, i.e., the 1-neighborhood [7, 35, 33], or as a parameterized version, the $d$-neighborhood structure [46]. When using this measure, the anonymization problem was shown to be NP-complete [33], making exact algorithms infeasible. Consequently, work has introduced greedy algorithms that modify node neighborhoods to resemble those of similar nodes [33]. Other methods include a genetic algorithm [35] and an edge sampling method [7].
@@
Graph invariants. The work of [26] introduced two novel measures based on graph invariants: 1) the number of nodes and edges in the $d$-neighborhood of a considered node, and 2) the degree distribution of the node’s $d$-neighborhood.
@@
VRQ. The notion of Vertex Refinement Queries (VRQ) was introduced in the work of Hay et al. [43, 22], accounting for the degree of nodes that are at most at distance $d$ of the considered node. To anonymize the networks, algorithms were introduced that use random rewiring [43] and alterations based on finding similar neighborhoods [47]. The work in [48] proves that anonymization using this measure is NP-complete for the subset variant of the problem, for which the aim is to anonymize a given subset of nodes.
@@
In this paper, we contribute to the existing literature by introducing and exploring three meaningful variants of the anonymization problem. We present a general framework applicable to all three variants of the anonymization problem, and contribute to existing methods in the literature by introducing several measure-agnostic anonymization algorithms. These can be used independently of the chosen anonymity measure. Additionally, we examine the impact of different anonymity measures on the anonymization process and analyze the trade-off between anonymity and utility.
@@ 3 Preliminaries
In this section we introduce concepts and terminology used throughout this paper. This includes general concepts for networks and anonymity, the alteration operations and the used anonymity measures.
3.1 Networks
We define a network as an undirected graph $G=(V,E)$ containing a set of nodes $V$ and a set of edges $\{v,w\}\in E$ between nodes $v,w\in V$. The degree of a node equals the number of connections it has: $degree(v)=|\{w:\{v,w\}\in E\}|$. The distribution of node degrees in a real-world network typically resembles a powerlaw distribution with many low degree nodes, and few high degree nodes [49].
Nodes in networks often tend to cluster. This tendency is captured for a node $v\in V$ by the clustering coefficient which equals the number of triangles it is part of, divided by the maximum number of triangles it could have: $\frac{1}{2}degree(v)(degree(v)-1)$. The clustering coefficient of a network equals the average clustering coefficient of all nodes in the network. The distance between two nodes, $dist(v,w)$, equals the number of edges that needs to be traversed to move from node $v$ to node $w$. It follows that $dist(v,v)=0$ and if there is no path between the nodes we define $dist(v,w)=\infty$. This occurs when two nodes are in a different component. The component in the network with the largest number of nodes is called the largest connected component (LCC), also referred to as the giant component.
We define the $d$-neighborhood of a node, $N_{d}(v)=(V_{N_{d}(v)},E_{N_{d}(v)})$, as the subgraph consisting of the set of nodes that are at most at distance $d$ of the considered node $v$, and the set of all edges between these nodes. Two $d$-neighborhoods are structurally indistinguishable if they are isomorphic, i.e, there exists a bijection between the two node sets that preserves all connections. We can determine whether two $d$-neighborhoods are isomorphic by comparing their canonical labeling [50]. This is a label assigned by a function $\mathcal{C}(G)$ such that two graphs have the same label value only if they are isomorphic.
Networks often consist of communities, being groups of nodes more strongly connected with each other, than with the rest of the network. Communities, i.e., clusters of tightly connected nodes, can be identified using community detection algorithms [5]. Such algorithms assign each node to a separate community, maximizing the number of intra-community connections and minimizing inter-community connections. To account for the randomness inherently present in these heuristic algorithms, resulting in different communities for the same graph, a consensus clustering step can be applied [51].
Nodes in a network play different roles, with some nodes being in a more central position than others. Various centrality measures have been introduced, which can be used to rank nodes [52]. A commonly used measure is betweenness centrality [53], which assigns a centrality score to each node based on the fraction of shortest paths that pass through this node, as denoted in Equation 1. Here $\sigma(s,t)$ equals the number of shortest paths from node $s$ to node $t$, and $\sigma(s,t|v)$ the number of such paths that pass through node $v$.
$c_{B}(v)=\sum_{u,w\in V}\frac{\sigma(u,w|v)}{\sigma(u,w)}$ (1)
@@ 3.2 Anonymity in networks
For a graph and an anonymity measure $M$, two nodes $v,w\in V$ are equivalent if $M(v)=M(w)$. Two such nodes are said to be in the same equivalence class, denoted $eq_{M}(v)$ for a node $v$. A node is $k$-anonymous if $|eq_{M}(v)|=k$. We say that $P_{M}$ denotes the partition of nodes into equivalence classes using measure $M$. When choosing $k=2$, the anonymity of a graph can be summarized using the metric of uniqueness. This is defined as the fraction of unique nodes in the network, as shown in Equation 3.
$V_{u}=\{v\in V:|eq_{M}(v)|=1\}$ (2)
$U(G)=\frac{|V_{u}|}{|V|}$ (3)
Additionally, we denote the set of unique edges, i.e., edges for which at least one of the nodes is unique, as follows.
$E_{u}=\{\{v,w\}\in E:v\in V_{u}\vee w\in V_{u}\}$ (4)
@@ 3.2.1 Alteration operations
When anonymizing a graph, the aim is to make as many nodes as possible $k$-anonymous, given a chosen measure for anonymity, by means of graph alterations. Applying these alterations results in an anonymized graph $G^{\prime}=(V^{\prime},E^{\prime})$. We experiment with the three edge alteration operations listed below. It should be noted that in many downstream tasks where the objective is to classify a particular aspect of a node, or to cluster nodes, operations that add or delete nodes are often not desired. Therefore, we do not consider node deletion or addition. Moreover, as preliminary experiments in Section 6.2 show that edge deletion is the most effective of the three, we focus on edge deletion in the remainder of the paper.
  • Deletion: $E^{\prime}\leftarrow E\setminus\{\{v,w\}\},\ s.t.\ \{v,w\}\in E$.
  • Addition: $E^{\prime}\leftarrow E\cup\{\{v,w\}\},\ s.t.\ \{v,w\}\notin E$.
  • Rewiring: $E^{\prime}\leftarrow E\setminus\{\{v,w\},\{v^{\prime},w^{\prime}\}\}\cup\{\{v,w^{\prime}\},\{v^{\prime},w\}\}$
@@
3.2.2 Measures for k-anonymity
Various measures for $k$-anonymity have been introduced in the literature. As described in [26], the choice of measure essentially determines the attacker scenario one protects against and has an effect on the uniqueness of the network. The measures both differ in what structural information is taken into account, and how far they reach. In this paper, we use the measures listed in Table 1. We choose one measure for each category in Section 2, aiming to represent the most common measures used in the literature. For completeness, the measures are parameterized by distance $d$, which indicates up to which distance structural information is taken into account when deriving equivalence classes. In this paper, we focus on $d=1$.
During the anonymization process we can assign a state to each node equal to the value computed by the anonymity measure, as denoted by column $M(v,d)$ in Table 1. The simplest measure, degree, is the degree of the node which can be determined in constant time. More complex are count and $d$-$k$-anonymity. To determine the value for a given node $v$ for these measures, the direct neighborhood of the node needs to be extracted. For count the number of nodes and edges needs to be extracted and for $d$-$k$-anonymity the canonical labeling of the neighborhood needs to be computed, which can be computationally expensive for larger neighborhoods. Lastly, the vrq measure accounts for the degrees of neighboring nodes. Compared to count and $d$-$k$-anonymity this measure is less expensive to compute, while it accounts for structural information reaching one edge beyond the neighborhood.
Table 1: Anonymity measures and their abbreviation (left two columns), definition given a graph $G=(V,E)$, considered node $v$ and distance $d$ (third column), reach type of the measure, and node set affected when altering edge $\{v,w\}$ (last two columns).
Measure Abbreviation $M(v,d)$ Measure type $A_{M}(\{v,w\})$